find the measure of an angle such the difference between the measure of its supplement and 3 times the measure of the complement is 16 degrees
Found 2 solutions by Gogonati, stanbon:
Answer by Gogonati(855)

Denote the unknown angle x degree, then its supplement is (180-x)degree and its complement is (90-x) degree. Since the difference between the supplement and three times the complement is 16 degree, we write the equation:
(180-x)-3(90-x)=16. The measure of unknown angle is 53 degree.

Answer by stanbon(75887)

find the measure of an angle such the difference between the measure of its supplement and 3 times the measure of the complement is 16 degrees
-----------
angle: x
complement: 90-x
supplement: 180-x
--------------------------
Equations:
180-x -(3(90-x)) = 16
180-x -270+3x = 16
2x - 90 = 16
2x = 106
x = 53 degrees
===================
Cheers,
Stan H.
=================